Displaying report 1-1 of 1.
Reports until 16:46, Friday 04 August 2023
Lasers General
Torrey Cullen - posted 16:46, Friday 04 August 2023 (11276)
More cryolab cavity locking

I moved the beam profiler to in reflection of the cavity to see if there was a discernable difference between the REFL beam when the cavity is locked and unlocked. There was not (see attached, should also more into this new beam profiler software and better ways to record data). I have not recovered alignment of the TRANS beam on the beam profiler. Aaron lent me a camera to use in transmission but we suspect the power incident on them is most likely too low for the poor QE camera. It seems like the TRANS beam should be much easier to find than this as the beam profile and camera diode are quite large. The cavity still appears to be locking on the mode we want.

To combat this I increased the power from the laser to 2.2mW coming from the source (allthough this puts you close to the maximum range on the TEC controller). This corresponds to ~1.3mW incident on the cavity. The 1811 PD saturates at far lower than this so I have a .5 ND filter in front of the PD.

Aaron thinks (from the elog I believe) that they were able to get up to 7mW out of the laser in the past. May be worth looking into this, but currently the range on the TEC controller doesn't allow this.

Images attached to this report
Displaying report 1-1 of 1.